OpenStudy (anonymous):

Really stuck on this one. Can someone walk me thru it?Which percent is equal to the given fraction? 1/20 A. 0.5% B. 5% C. 10% D. 20%

OpenStudy (bibby):

convert the fraction to a decimal, multiply by 100 to find the percent

OpenStudy (anonymous):

The given fraction is 1/20

OpenStudy (bibby):

do you know how to convert that to decimal using division?

OpenStudy (bibby):

alternatively you can just get it in terms of 100 like so \(\dfrac{1}{20}=\dfrac{x}{100}\)

OpenStudy (anonymous):

1/20=5/100?

OpenStudy (bibby):

exactly, and percent just means out of 100

OpenStudy (anonymous):

so. ugh. a?

OpenStudy (bibby):

b, 5/100 = 5% = 0.05 if percent is out of 100, and we have 5 out of 100, that means we have 5 percent

OpenStudy (anonymous):

Ohhh. Whoops. Didnt read it correctly. So I think i get it now.

OpenStudy (bibby):

do you want me to write out the steps more in depth?

OpenStudy (anonymous):

Uhh. I dont think so i think i got it. My next one is kinda tricky tho. I have to find a percent that is equivalent to 7 1/4 do i just do the same thing?

OpenStudy (bibby):

first get it into improper form

OpenStudy (anonymous):

Okay so to do that i 7x4+1?

hero (hero):

You don't need to convert Anything to improper form.

hero (hero):

Converting to improper form is not a correct approach.

OpenStudy (bibby):

if we're making the denominator 100, it is or she can just convert it to decimal there is no single correct approach, your input isn't appreciated if you aren't going to be constructive

OpenStudy (anonymous):

Now im confuzzled.

OpenStudy (bibby):

just do what I say until he can justify himself

OpenStudy (anonymous):

Okay.

OpenStudy (bibby):

also yeah, sorry I didn't see that post earlier. \(7 \dfrac{1}{4} = \dfrac{7*4+1}{4} =\dfrac{29}{4}\) then we want \(\dfrac{29}{4}=\dfrac{x}{100}\) alternatively, we can just convert it to decimal using a calculator. \(7\dfrac{1}{4}=7.25\) then we divide by 100 to find the percent. \(\dfrac{7.25}{100}=x%\)

OpenStudy (bibby):

\(\dfrac{7.25}{100}=x\%\) *

OpenStudy (anonymous):

uhh.

OpenStudy (bibby):

dividing by 100 means we move the decimal two places to the left

OpenStudy (bibby):

the basic rule is converting decimal to percent: move the decimal point two places right converting percent to decimal: move the decimal two places left